OpenStudy (anonymous):

Simplify, Assume that no variable equals 0: b^8/b^2

OpenStudy (anonymous):

If you're just learning powers I like to start by expanding the power out \[\frac{b^8}{b^2} = \frac{b*b*b*b*b*b*b*b}{b*b}\] So we can cancel 2 b's on the top and bottom and we're left with.. \[b*b*b*b*b*b = b^6\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

a lil bit better but that is only one of the simpler problems on my worksheet

OpenStudy (anonymous):

Well I think you'll find that a similar approach works well on even more complicated problems. But if you have trouble feel free to ask.